1. log6 + Log3 - Log X = Log9
2. log 125
3. log t
4. log 1

Answer by ilana(305)

For #1, we can condense this using log properties to log((6*3)/X)=log9.
We can then take 10^(each side), giving the algebraic equation 18/X=9, so 9X=18, so X=2.
#2 means 10^X=125.
#3 means 10^X=t.
#4 means 10^X=1, so X=log1=0.
I'm not completely sure what the question is asking of you in 2,3, or 4.